Results from the ATLAS Semiconductor Tracker Barrel System Test

P.J. Bell^{a}, P. Brückman^{b}, S. Gadomski^{c,1}, J. Grosse-Knetter^{c,∗}, M. Morrissey^{d,2}, K. Nagai^{e}, J. Pater^{f}, B. Skubic^{g,3}

aUniversity of Birmingham, Birmingham, UK bInstitute of Nuclear Physics, Cracow, Poland cCERN, Geneva, Switzerland dRutherford Appleton Laboratory, Didcot, UK eQueen Mary, University of London, London, UK fUniversity of Manchester, Manchester, UK gKungl. Tekniska Högskolan, Stockholm, Sweden

Abstract

Results are presented of tests with a partial prototype of the barrel part of the ATLAS SCT. Up to 17 non-irradiated and 5 irradiated modules have been tested on this system and showed a performance well within specifications. Different grounding and shielding configurations have been tried and sufficient protection from external noise has been achieved.

Key words: silicon detector, LHC, system test PACS: 06.60.Mr, 29.40.Gx

1 Introduction

ATLAS is one of the two multipurpose detectors for new physics in

∗ Corresponding author.

Email address:

joern.grosse-knetter@cern.ch (J.

Grosse-Knetter). 1 Now at University of Bern, Switzerland. 2 Now at University of Oxford, UK. 3 Now at XXX, Sweden?

the TeV-mass range which are constructed for operation at the LHC [1]. The ATLAS apparatus consists of several components such as superconducting magnets, muon chambers, calorimeters, and the Inner Detector (ID). The ID is designed for precision tracking of charged particles with 40 MHz bunch crossing identification [2]. It combines tracking straw tubes in the outer transition-radiation tracker (TRT) with semiconductor microstrip de-

Preprint submitted to Nuclear Instruments and Methods 26 April 2002

tectors of the Semiconductor Tracker (SCT) in the middle and semiconductor pixel detectors in the innermost PIXEL component.

The SCT [2–4] consists of four nested barrels in the centre and nine disks each forming the end-caps. The barrels cover a region from 300 to 520 mm in radius and have an active length of 1530 mm, centred around the interaction point. The respective barrel layers are fully covered by 32, 40, 48, and 56 rows of twelve identical modules, ie a total of 2112 modules.

The SCT modules [4,5] are made of four sensors, glued in pairs on either side of a thermally highly conductive baseboard. The sensors are approximately 6×6 mm^{2} p-in-n silicon wafers of approximately 300 μm thickness. Each sensor has 768 strips with 80 μm pitch. The modules are equipped with hybrids holding twelve read-out chips, six on the top and six on the bottom side of the module. Such a chip provides highgain (50mV/fC) fast-shaping (20ns peaking time) binary read-out of 128 channels in radiation-hard technology. All channels have a common threshold for discrimination, which can be adjusted individually for each channel by a 4-bit trim DAC (trimming) to cope with the increased offset spread after irradiation. The heat produced by the chips and the detectors is extracted via the baseboard at one end of the module. Tests with individual prototype modules showed a performance well within the specifications [5,6].

The goal of the system test is to run as many modules as possible in a physical configuration which is as close as possible to the planned ATLAS SCT configuration, thereby testing the performance of the modules in such a system and comparing to their stand-alone performance [3,5,6]. The system test complements beam tests and irradiation tests [6–10] in assuring full functionality of the final SCT system.

2 Set-up of the system test

The system test simulates a section of one of the SCT barrels in all its details of mounting, cooling, supply, and shielding. Figure 1 shows a schematic drawing of an SCT barrel in the design as used for the system test set-up^{4} and figure 2 shows a photograph of the system test setup. Specifications of the location of objects on the sector (left, front, etc) refer to the orientation as shown in figure 2.

2.1 Mechanical support

Modules are mounted on a sector of a carbon-fibre-Corex-sandwich cylinder with dimensions very near to those of the innermost ATLAS SCT barrel. The mechanical support of

4 Several details in the design have changed since construction of the system test set-up, such as the shape of the cooling pipes, but are thought to have no or negligible relevance to the performance tests discussed here.

Fig. 1. Drawing of a section of a SCT barrel in 3D-view in the design as used for the system test set-up.

Fig. 2. Photograph of the barrel system test showing a configuration with one row completely equipped with modules.

modules is realised with carbon-fibre brackets which are mounted on the Corex sector. The sector can hold up to 48 modules, in four rows of twelve. The rear row is not equipped, as are the right halves of the two front rows.

Above sector is mounted on three aluminium frames which are supported by a carbon-fibre rod. The rod itself is fixed on either end by vertically placed aluminium plates. The whole system sits on an aluminium support structure.

A plastic cover with square crosssection, complementing the aluminium support, is used to provide a gas- and light-tight enclosure. The enclosed volume is flushed with dry nitrogen while running thus providing an atmosphere with a relative humidity of 20% or less.

2.2 Cabling

Modules are powered and read out via prototype opto-harnesses [11], which transmit low and high voltage to the modules. Data and command transfer employs optical signals (see section 2.4 below) and these signals are processed with dedicated circuits on the harnesses [12]. They decode the incoming optical clock and commands, transmit these signals electrically to the modules, and transform the data from the modules into optical signals for transfer to the read-out electronics. Each optoharness serves up to six modules. Each harness consists of six 50 μmaluminium-on-kapton tapes and six bunches of three optical fibres. Six doglegs are interfacing the power lines on the main tapes to each module and carry the optical circuits. Thus two harnesses are needed per row of twelve modules. These two harnesses are separated from each other and connections are made only at the respective ends of the barrel sector.

All patch panels and power tapes used are true to the planned final ATLAS design, except for extra provisions on the patch panels to allow testing of various coupling schemes.

The above is actually not quite

                                 true any more for PP2, PP3. The

opto-harness ends are connected to a first patch-panel (PP1) that provides filtering on the power lines (see section 2.6 below). It interfaces to 4 m of 100 μm-aluminium-on-kapton power tapes which connect to the next patch panel (PP2). The following section is not very close to the ATLAS design and a simplified version of the foreseen third patch panel (PP3) is directly connected to PP2. The conventional cables used between PP3 and the power supplies are 30 m long and thus shorter than what is needed for ATLAS, however the cable resistance is kept similar. The optical signals are transmitted through 25 m long fibres.

2.3 Cooling

The hybrids and detectors of each module produce between 5 and 10 W of heat which needs to be taken away. For this purpose the three front rows on the sector are equipped with aluminium pipes. They have a flat-oval cross section onto which the module baseboards are clamped. The pipes are mounted into an aluminium support bracket attached to the sector on either side. An ethanolwater mixture is pumped through the pipes with the temperature being controlled by a commercial chiller.

2.4 Read-out

The modules are read out using a VME-based prototype system [13]

Fig. 3. Schematic diagram of the system test DAQ-system.

that consist of four parts as shown in figure 3. Two parts, Cloac and Slog, generate clock, command and trigger signal, another, Mustard, decodes the data received from the modules. All three are connected to an opto-transceiver, Optif, to transform clock and command from electrical into optical, and the received data from optical into electrical signals. It connects to the fibres coming from the harnesses. These DAQ components are used in place of the finaldesign read-out drivers (ROD) [2]. The low-voltage power supplies (see section 2.5 below) are also used to read module temperatures for monitoring purposes. A ROOT-based DAQ-software package is used [14], running on a Windows-NT PC which is connected to the VME crates via a National Instruments interface card.

2.5 Power supplies

Power for the chips on the modules and the optical circuits on the doglegs is provided by prototype VME supplies (SCTLV); detectors are currently biased with the companion SCTHV units [15]. These units provide two separate low-voltage channels (one for the analogue circuits and one for the digital circuits on the modules, later referred to as analogue/digital voltage/ground) and one separate high-voltage channel for each module. The individual channels are floating with respect to each other. It should be noted, however, that digital, analogue, and high-voltage grounds are connected on the modules.

2.6 Shielding

The system test bases its default grounding and shielding scheme on the proposal for the final detector [16,17]. The main elements as applied in the system test are described below, with any differences noted.

To control stray capacitance between the cooling pipe and the SCT detector backplane, shunt shields are placed between the modules and the cooling pipe. These shunt shields are 50 μm(?)-copper-on-kapton pads of 13×75 mm^{2} in size; the copper is soldered to analogue ground on the dogleg.

The cooling pipes are connected together at both ends of the sector to their support bracket. This bracket is connected to the aluminium parts of the support structure which provide the reference point of the shield. The pipes are split in the centre where they are glued into a plastic-joint providing electrical insulation.

The sector is also connected electrically to the metalwork which supports the sector. A cover made of fibre-glass with a 100 μm thick aluminium foil on its underside is placed approximately 10 cm above the carbon fibre structure and simulates the thermal shield. It is connected to the main metal structure on either side.

The six power tapes for each harness are wrapped together with 50 μmaluminium-on-kapton shielding from the end of the thermal shield to PP1. The wrap is DC-connected to the metalwork surrounding the sector. At PP1 digital and analogue grounds are AC-connected (with 2.7 μF) to the cable shielding - Ned wants a

DCconnectionbetweenshieldand

Dgnd. PP1 provides similar capacitive filtering of all other power lines to the cable shielding. The shield on the PP1s on either side of the sector are connected to two 300×900 mm^{2}- sheets of 100 μm thick copper, respectively, which are connected to the main metal structure with their other end. The sheets are used to simulate the extra shield connection provided by the radial heat spreader plated running radially outwards from the thermal shield to the PP1 location in the final SCT detector.

From PP2 to PP1 the system test shield is not very like the ATLAS SCT; there is no surrounding metal (e.g. heat spreader plates, cooling supply tubes) to connect to as is recommended [16]. However, the power tapes are wrapped in bundles of six in the same aluminium-on-kapton foil as the harness ends. The wrap is connected to the respective copper sheet at PP1.

Capacitive filtering to the shield and common-mode chokes are used at PP3 but as discussed above there is no extra cable between PP2 and PP3. The conventional cable shield is connected through at PP3 and connects to the tape wrap at PP2 and to the respective digital grounds on the low-voltage supplies.

An alternative grounding and shielding scheme was tested [17,18]. It involves DC connections between all metal in the system and DC connection of the module grounds to the cooling pipe. The latter is realised by shorting the shunt shields used in the default set-up. In addition, electrically continuous cooling pipes are used and the cable shields are ACconnected to VME metal instead of digital ground on the low-voltage supplies.

2.7 Cold operation

In order to simulate the situation of running at low temperatures in ATLAS a cryophasing procedure has been established to safely cool down the whole sector to temperatures below zero, take data, and subsequently warm it up again. One important safety issue is the possibility that condensation may form on sensitive components on the sector (modules, opto-chips) so that during the whole cryophasing the dew-point is measured by a slow-control system and used as a safety-limit. The actual cool-down or warm-up is performed using the chiller normally only cooling the pipes together with a commercial freezer.

Cold operation is a prerequisite for the operation of irradiated modules on the sector due to the possibility of thermal runaway caused by the increased leakage currents drawn by irradiated detectors.

Operating conditions after cool-down correspond to an ambient temperature of -10^{∘}C and a pipe temperature of -15^{∘}C.

3 Analysis methods

This section describes different methods of testing the noise performance of the system. These methods use one of the following two data taking modes:

• random triggers; • random triggers with calibration

pulses: the calibration circuits on the module hybrids are used to inject a well defined charge into the preamplifiers, individually for each detector strip.

In both cases the data taking is sequential, ie one trigger is sent, the data is read and decoded, after which another trigger is issued etc.

3.1 Calibration scans

This type of measurement allows to determine the amplifier gain and the noise charge on the preamplifier input for each channel on a module [14]. The occupancy is measured with random triggers plus calibration pulse while scanning versus amplifier threshold. The response curves are fitted with an error function, yielding a relation between injected charge and threshold at 50% occupancy so that several scans with different injected charge allow to determine the gain. The Gaussian σ of the error function is the noise as seen on the amplifier output and is converted to input-noise charge Q_{noise} using the gain.

The noise values quoted below are determined with a gain-measurement at injected charge of 1.5, 2, 2.5 fC and the actual noise-value is taken from the 2 fC scan.

3.2 Noise occupancy scan

The noise occupancy is measured scanning versus threshold charge, Q_{thr}, with the number of random triggers being varied between 2000 and 10^{6} to adjust for decreasing number of hits as the threshold increases. The calibration of the Q_{thr}-scale is as described in section 3.1.

In good approximation the logarithm of the occupancy O depends linearly on the square of the threshold for Q_{thr} ≫ Q_{noise} [19]:

ln(O) ≈ −\frac{1}{2} \frac{Q_{thr}^{2}}{Q_{noise}^{2}} . (1)

Deviations from the linear behaviour indicates (non-Gaussian) noise pickup. If a linear behaviour is observed an according fit for Q_{thr} > 0.5 fC is used to determine Q_{noise} which is typically around 0.2 fC.

This method offers a way to measure Q_{noise} alternatively to above calibration scan. It should be noted, however, that such a calibration is needed in order to determine Q_{thr} so that this method and the calibration scan can not be seen as totally independent of each other.

3.3 Repeated noise measurements

The calibration measurement described in section 3.1 is repeated over several hours. Both the variation of the gain and of the noise is analysed. For many modules in operation this measurement is very slow – it typically takes about 1/2-hour per individual measurement.

In a different mode, the noise occupancy is measured repeatedly in a random trigger mode. Repeated bursts of 10^{6} triggers at a threshold of Q_{thr} = 1 fC are used. Both the occupancy as a function of time and the average occupancy are analysed. This method is much faster than the above, it typically takes a few minutes for an individual measurement, which makes it very sensitive to fluctuating noise pick-up.

3.4 Common mode noise

Common mode noise pick-up is a particularly worrying problem for binary read-out systems as used for the SCT modules. Thus special analysis methods are dedicated to this problem [19]. Data is taken in a random triggering mode, storing the full hit information for each event. This information is then used to calculate a correlation matrix or compare the spread of the number of hits on a chip or module to the theoretically expected value without pick-up (Γmethod). Either method allows to determine the magnitude of potential pick-up, without need of comparison to “bench mark” measurements taken outside the sector [20].

3.5 Noise injection

In the final SCT the thermal shield and other metal parts running along the barrel axis short the cable wraps and the cooling supply-pipes coming in from either side of the barrel, very much like in the system test. Therefore significant noise current might flow through such parts of the shield.

To enforce such a situation, a sinusodial noise current is injected into the thermal shield and the noise is measured with a calibration scan (see section 3.1). The injected current comes from a commercial signal generator of which the ground is connected to the main metal work to one side of the sector. The signal output is connected via a 1 Ω resistor to the other end of the thermal shield. The resistor is used to measure the current that flows, but it should be noted that the current is divided between the thermal shield and the main metal work due to the shielding connections made between both. The noise current is quoted as peak-to-peak value. The noise measured with a calibration scan with noise injection is compared to a measurement with the signal generator still connected but set to zero output.

3.6 Crosstalk

                                 Ifweevergetthisgoing...

4 Individual tests

When a module arrives at the system test, it is accompanied by the results from a standard characterisation as performed at the manufacturing institute [14]. The first step in integrating the module into the system test is to repeat this standard characterisation on an electrical test bench to verify that the module has not suffered in transit. The electrical test bench is considerably simpler than the full system test as it bypasses the optical communication, employs only very short power and signal cables, and has well controlled grounding and shielding since the module is mounted in a metallic box. Therefore a module running standalone on the test bench tends to give its best performance. The noise figures from the calibration scans are used as bench marks when analysing similar data from multi-module runs on the sector.

Once the module is verified to be in good working order, it is mounted on the system test sector with all grounding and shielding connections made, and the standard characterisation sequence is repeated, powering only that one module, and this performance is compared to its performance on the bench. Any differences are noted, and investigated if possible. When this comparison is complete, the module is considered ready to be included in multi-module

tests. Shall we add a list of all
ourmoduleshere?

Part of the characterisation sequence applied during these individual test are a full trimming and a calibration run [14]. The results of these are used to initialise and calibrate all modules in all subsequent multi-module runs.

5 Tests with several modules

                                 Thefollowinglistsmeasurements
                                 I  would  like  to  see  in  this
                                 section. Those marked with (?)
                                 should probably not be in this
                                 paper (in particular those in
                                 5.4), but those are at least
                                 testswedid

5.1 Tests in default set-up

                                 3-ptgainresults+comp.toelec
                                 stand;
                                  averageoccupancyat1fC+comp.

Fig. 4. ENC-noise multi-module vs electrical stand.

                                 toelecstand;
                                  stability(gain,noise,occupancy);
                                 noiseoccupancyvsQthrscans;
                                 correlatednoiseresults;
                                 noiseinjection;
                                 crosstalk(ifwegetthisdone);
                                 row-of-12 -- full+half-row --
                                 3-half-rows(ifwehave18modules).

Fig. 5. Average noise occupancy: multi-module vs electrical stand.

5.2 Shielding studies

Ned vs Tony (ie shunt shields
in/out,(un-)splitpipes);
other "optional" parts in the
shield like dummy-radial heat
spreaderplates,PP1/chokefiltering(?);
effectofbadconnections(?).

5.3 Cold running and irradiated system test

asinthetitle

5.4 Misc.

redundancy links in operation
(requires extra explanation in
2.2)(?);
testswithHeinz'PP1-PP2cables(?);
LV2-LV3comparison(?);
stabilitydegradationwithreduced
biasvoltage(?).

6 Summary

All’s well that ends well.

Acknowledgements

The efforts of R. Fortin and B. Herskind as well as G. Barbier and E. Perrin during the production of components for the system test and their maintenance are highly appreciated. The helpful and competent assistance of P.W. Phillips with all our software and DAQ-related problems was invaluable. We acknowledge the advice of N. Spencer and T. Smith in setting up and improving a grounding and shielding scheme. Shall we

                                 mentioninstitutedonations?

References

[1] ATLAS Technical Proposal for a General-Purpose pp Experiment at the LHC at CERN, CERN/LHCC/94-43 (1994).

[2] ATLAS Collaboration, ATLAS Technical Design Report - Inner Detector, CERN/LHCC/97-16 and CERN/LHCC/97-17 (1997).

[3] M. Turala for the ATLAS SCT collaboration, Nuclear Instr. Meth. A 466 (2001) 243.

[4] Y. Unno for the ATLAS SCT collaboration Nuclear Instr. Meth. A 453 (2000) 109.

[5] T.  Kondo  for

the ATLAS SCT collaboration, 5th International Conference on Large Scale Applications and Radiation Hardness of Semiconductor Detectors, Firenze, Italy (2001).

[6] H. Pernegger for the ATLAS SCT collaboration, VERTEX2001, Brunnen, Switzerland (2001).

[7] Y. Unno for the ATLAS SCT collaboration, 2001 IEEE NSS/MIC, San Diego, USA (2001).

[8] M.  Vos  for

the ATLAS SCT collaboration, 5th International Conference on Large Scale Applications and Radiation Hardness of Semiconductor Detectors, Firenze, Italy (2001).

[9] A. Barr et al, ATLAS note ATL-INDET-2002-005 (2002).

[10] R.S. Harper et al, ATLAS note ATL-INDET-2000-009 (2000).

[11] T.  Weidberg,  ATLAS-

EDMS technical note ATL-IS-ES-0049 (2001).

[12] D.G. Charlton et al., Nuclear Instr. Meth. A 443 (2000) 430.

[13] M. Morrissey, M. Postranecky, M. Warren, et al., documentation of the Slog, Cloac, Mustard DAQ system, available at: http://hepwww.rl.ac.uk/atlassct/mm/Slog/slog.pdf http://www.hep.ucl.ac.uk/atlas/ sct/cloac/Welcome.html http://hepwww.rl.ac.uk/atlassct/mm/Mustard/Mustard.ps

[14] P.W.  Phillips,  ATLAS-

EDMS technical note ATL-IS-??? (not out yet, called Operation of Barrel Modules).

[15] P. Malecki for the ATLAS SCT Collaboration, 7th Workshop on Electronics for LHC Experiments, Stockholm, Sweden (2001).

[16] N.  Spencer,  ATLAS-

EDMS technical note ATL-IC-EN-0004 (1999).

[17] M. Morrissey and T. Weidberg, ATLAS-EDMS technical note ATL-IS-ES-0056 (2001).

[18] T. Smith, ATLAS-EDMS technical note ATL-IS-ER-0019 (2001).

[19] L. Feld, S. Roe, A. Ahmad, S. Snow, ATLAS note ATL-INDET-2001-006, to appear in Nucl. Instr. Meth.

[20] P.J. Bell et al., our common mode noise write-up.